OpenStudy (anonymous):

how do i solve: x=sin(-16\pi)

OpenStudy (anonymous):

if you want \[\sin(\frac{16}{\pi})\] you are going to need a calculator

OpenStudy (anonymous):

what if i dont have a calcultor though.... like there's no other way of solving that?

OpenStudy (anonymous):

same for \[\sin(-\frac{16}{\pi})\] you are not going to find it any other way

OpenStudy (matlee):

Nope its impossible because pie is continuous

OpenStudy (anonymous):

no there isn't and of course you have a calculator, you are on a computer my guess is that the question really says something else, but i could be wrong

jimthompson5910 (jim_thompson5910):

If you meant to say \[\Large x = \sin(-16\pi)\] then use the rule sin(n*pi) = 0 n is any integer

OpenStudy (anonymous):

alright, thank you

OpenStudy (anonymous):

\(\pi\) is not "continuous" \(pi\) is a number can you post a sceenshot of the actual quesion?

jimthompson5910 (jim_thompson5910):

as for a calculator, you can use wolfram alpha

jimthompson5910 (jim_thompson5910):

http://www.wolframalpha.com/

OpenStudy (matlee):

pi is not continuous? there is only an esitmate

OpenStudy (anonymous):

x=sin(-16pi)

OpenStudy (anonymous):

that's what the question is

OpenStudy (matlee):

did u get the answer

OpenStudy (anonymous):

nope, it doesn't make sense

OpenStudy (matlee):

what did you get. i got 0

OpenStudy (anonymous):

wait yeah i got 0 but is that all i do? put it in a calculator?

OpenStudy (matlee):

To tell you the truth i would do the same thing you are but yeah its that easy -_-

OpenStudy (matlee):

Let me try to find the long way

OpenStudy (matlee):

You can learn how to use sine this way http://mathworld.wolfram.com/Sine.html its under the number 3

OpenStudy (matlee):

then pretty much multiply -16*pi and and then solve the rest of the equation